OpenStudy (anonymous):

Will give medalsss Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = quantity x minus nine divided by quantity x plus five. and g(x) = quantity negative five x minus nine divided by quantity x minus one.

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{x-9}{x+5}\qquad g(x)=\cfrac{-5x-9}{x-1} ?\)

OpenStudy (jdoe0001):

you can always post a screenshot here btw, just by using the [Attach File] blue button, so you know

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok... so if indeed, f(x) and g(x) are inverses of each other that means that g(x) is the inverse of f(x) and that f(x) is therefore the inverse of g(x) so to verify, you'd do as indicated find f( g(x) ) and that will simplify to "x" if they're indeed inverses of each other lemme do the first one, so you can do the next one

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{x-9}{x+5}\qquad {\color{brown}{ g(x)}}=\cfrac{-5x-9}{x-1} \\ \quad \\ f(\quad {\color{brown}{ g(x)}}\quad )=\cfrac{{\color{brown}{ g(x)}}-9}{{\color{brown}{ g(x)}}+5}\to \cfrac{\left({\color{brown}{ \frac{-5x-9}{x-1}}}\right)-9}{\left({\color{brown}{ \frac{-5x-9}{x-1}}}\right)+5}\to \cfrac{\frac{-5x-9-9(x-1)}{x-1}}{\frac{-5x-9+5(x-1)}{x-1}} \\ \quad \\ \to \cfrac{\frac{-5x\cancel{ -9 }-9x\cancel{ +9 }}{x-1}}{\frac{\cancel{ -5x }-9\cancel{ +5x }-5}{x-1}}\to \cfrac{\frac{-14x}{x-1}}{\frac{-14}{x-1}}\to \cfrac{\cancel{ -14 }x}{\cancel{ x-1 }}\cdot \cfrac{\cancel{ x-1 }}{\cancel{ -14 }}\) notice what's left so now... try the 2nd one g( f(x) )